# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

### 29 Feb 2024 v1

@@ 7. Conclusion
This paper introduced a novel graph representation of the programming language as a complex-valued multigraph with a polynomial-based adjacency matrix encoding. It evaluated its applicability on the declarative specification language Alloy. It further evaluated its improvement of compactness on a dataset consisting of multiple Alloy models and the comparability of the Alloy predicates in an attempt to develop an automatic fixing framework. It points out a direction of inquiry in utilizing the representation in multiple software engineering practices utilizing Alloy and possible integration in a machine learning algorithm.

### 21 Mar 2024 v3

@@ 7. Conclusion
This paper introduced a novel graph representation of the programming language as a complex-valued multigraph with a polynomial-based adjacency matrix encoding. It evaluated its applicability on the declarative specification language Alloy. It further evaluated its compactness improvement on a dataset consisting of multiple Alloy models and the comparability of the Alloy predicates in an attempt to develop an automatic fixing framework. It points out a direction of inquiry in utilizing the representation in multiple software engineering practices utilizing Alloy and possible integration in a machine learning algorithm.